OpenStudy (anonymous):

what is the inverse of y = 3x² - 11x - 4

OpenStudy (johnweldon1993):

switch all the x's and y's y = 3x² - 11x - 4 becomes x = 3y² - 11y - 4 now solve for y again

OpenStudy (anonymous):

what would you do after \[x+4=3y ^{2}+11y\]

OpenStudy (anonymous):

I meant to say x+4=3y^2-11y

OpenStudy (anonymous):

from: x = 3y² - 11y - 4 x+4 = 3y² - 11y - 4 x+4 = y ( 3y - 11) \[y=\frac{x+4}{3y-11}\]

OpenStudy (anonymous):

the inverse of y=mx+c = y ^ -1 (that is inverse, swaping x n y) therefore it would be x=my+c, solve for x therefore 1. x - c = my 2. (x - c) / m = y, hope this helps

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The inverse of a squared expression is going to be an expression involving a square root.